Analyze the two following situations for firms in competitive markets:
a. Suppose that TC = 100 + 15q, where TC is total cost and q is the quantity produced.
What is the minimum price necessary for this firm to produce any output in the short
run?
b. Suppose that MC = 4q, where MC is marginal cost. The perfectly competitive firm
maximizes profits by producing 10 units of output. At what price does it sell these
units?
1
Expert's answer
2018-04-18T09:53:08-0400
Analyze the two following situations for firms in competitive markets:
a. If TC = 100 + 15q, then the minimum price necessary for this firm to produce any output in the short run is P = AVC, AVC = (TC - FC)/q, so P = AVC = 15q/q = 15.
b. If MC = 4q and the perfectly competitive firm maximizes profits by producing 10 units of output, then the profit maximizing price is P = MR = MC = 4q = 4*10 = 40.